# Buy versus Build an LLM: A Decision Framework for Governments

### v1

@@ 4.5 Resource Capability and Sustainability Perspective
  • Talent Base. A successful build strategy requires a strong team of experts across hardware infrastructure, data engineering, algorithm, security and governance. The central challenge lies in the government’s ability to compete with the private tech companies to attract and keep these professionals, who can often be recruited by tech firms with high-paying job offers. It is obvious that having a strong national research-intensive universities is critical for attraction, retention and sustained production of talent. Some countries rely on international talent while that might not be an option for other countries.
  • Hardware, Infrastructure and Utilities A sovereign build strategy requires a robust physical foundation, including high-performance GPUs, low-latency networking and a massive, reliable supply of electricity and cooling. A build strategy is only viable if the country has reliable access to these physical assets or maintains a strategic reserve to withstand global market volatility.
  • Environmental Sustainability. Large-scale language models can be energy-intensive across training, deployment, and ongoing operations, making environmental sustainability a critical consideration. Building and operating an in-house model entails a non-trivial electricity demand and associated carbon footprint, which may conflict with national climate commitments unless paired with low-carbon power procurement and efficiency measures. For instance, the training of the 176176B BLOOM model consumed $433,196$ kWh total energy and was estimated to emit about $24.69$ tCO2[57]***tCO2e: (tonnes of carbon dioxide equivalent), a standard measure that expresses the total climate impact of greenhouse gas emissions on a common CO2 basis.. Importantly, environmental impacts do not end at training: the inference stage can accumulate substantial footprint at scale. Mistral’s lifecycle assessment [95] reports that by January 2025, Mistral Large 2’s training plus 18 months of usage amounted to $20.4$ ktCO2e, $281,000$ m3 of water consumption, and $660$ kg Sb eqkg Sb eq: (Kilograms of antimony equivalent), a life-cycle assessment metric indicating depletion of scarce mineral resources, normalized to antimony as a reference. of resource depletion. On average, a single $400$-token response (excluding end-user devices) contributes an estimated $1.14$ gCO2e, $45$ mL of water, and $0.16$ mg Sb eq [95]. Under a buying strategy, these impacts are often operationally externalized to the provider’s infrastructure, but governments may still face indirect accountability through procurement standards, public reporting expectations, and alignment with national sustainability targets.
@@ 4.6 National Context Fit Perspective
  • Language. Language coverage is one of the core determinants of the decision especially in multilingual societies. If the primary language is already well supported by global tech companies, buying can deliver strong performance with minimal upfront effort. In contrast, for the purpose of serving minority languages and regional dialects where commercial models may struggle with, the building strategy allows governments to prioritize these low-resource languages using their own specialized datasets. Building can therefore improve inclusive access to public services and help preserve linguistic identity, but it requires sustained effort in data collection, evaluation, and ongoing maintenance.
  • Cultural Fit. The ability to handle historical references, political context, and local social norms is vital for maintaining public trust. Cultural fit involves more than just language; it requires the model to respect religious beliefs, understand local taboos, know local cuisines and follow regional customs. Furthermore, many countries require fine-grained control to account for different laws or values across various states and provinces. In some countries, rich cultural traditions are gradually eroding as younger generations increasingly adopt global cultural norms. Governments may view culturally aligned LLMs as a tool to support cultural continuity, particularly in education, public communication and youth engagement. A locally developed or highly customized model can be programmed with these specific sensitivities in mind, ensuring that its outputs are not only accurate but also culturally appropriate and socially responsible.
  • Legal Interpretations. Can LLMs embed national legal frameworks, case law standards, and procedural rules into its parameters? An LLM should be able to embed national statutes, case law standards, and administrative rules directly into its parameters or retrieval systems. In practice, general-purpose models could default to assumptions from dominant jurisdictions from foreign laws, producing answers that may be misleading, inappropriate, or even unlawful in other countries. When pursuing a buying strategy, governments may therefore need to explicitly negotiate legal disclaimers and domain constraints with model providers to mitigate these risks. On the other hand, building carries risk: without sufficient data and building capability, such systems can produce misleading, hallucinated outputs or even illegal advice [105], wasting resources and undermining trust. A recent example is New York City’s internally developed legal chatbot, which was shut down after producing unreliable advice despite costing over $\mathdollar 600{,}000$ to build [105]. In such cases, combining reliable priopriety models with controlled retrieval (e.g., RAG) may be the safer choice.
@@ 4.7 Evolving Cost-Capability Landscape Perspective
  • Temporal Cost Dynamics. The buy-build decision can shift substantially over time as the underlying cost–capability frontier moves. Training costs may decline rapidly; for example, DeepSeek V3 reports a training cost of roughly \$5.6M, an order-of-magnitude reduction relative to earlier frontier-scale efforts. In parallel, the open-source ecosystem continues to mature, with open models increasingly approaching frontier performance. Compute hardware is also becoming more widely available, while local talent pools can deepen through knowledge transfer and the growing transparency of technical reports and academic publications. Together, these potential trends can reduce the cost and raise the feasibility of building, such that a government may rationally choose buy now and build in the next few years, not because its preferences changed, but because the underlying cost–capability landscape shifted.
  • Capability Debt. LLM development capability is not a function of capital investment alone; it is cumulative, depending on hands-on training experience, institutional know-how, data pipelines and partnerships, and operational lessons learned through iteration and failure. Delaying capability building can therefore create capability debt, raising the cost and difficulty of catching up even as absolute compute costs decline.
  • Option Value of Modest Early Investment. Modest early investments such as small-scale training pilots, sovereign benchmarks, and infrastructure prototyping can preserve future build options without immediate commitment to full-scale training. These efforts accumulate organizational know-how, reduce future transition costs, and enable faster scaling if priorities or constraints change. This option value is often omitted in static cost comparisons but can be strategically material.
  • Re-evaluation Triggers. Governments can make regular re-evaluation by setting clear triggers, such as widening capability gaps, major improvements in open models, vendor price or policy changes, new regulations, or clear progress in domestic talent and infrastructure. Defined triggers and review cycles ensure the strategy keeps pace with the technology landscape rather than changing only in response to crises.

### v2

@@ 4.5 Resource Capability and Sustainability Perspective
  • Talent Base: A successful build strategy requires a strong team of experts across hardware infrastructure, data engineering, algorithm, security and governance. The central challenge lies in the government’s ability to compete with the private tech companies to attract and keep these professionals, who can often be recruited by tech firms with high-paying job offers. It is obvious that having a strong national research-intensive universities is critical for attraction, retention and sustained production of talent. Some countries rely on international talent while that might not be an option for other countries.
  • Hardware, Infrastructure and Utilities: A sovereign build strategy requires a robust physical foundation, including high-performance GPUs, low-latency networking and a massive, reliable supply of electricity and cooling. A build strategy is only viable if the country has reliable access to these physical assets or maintains a strategic reserve to withstand global market volatility.
  • Environmental Sustainability: Large-scale language models can be energy-intensive across training, deployment, and ongoing operations, making environmental sustainability a critical consideration. Building and operating an in-house model entails a non-trivial electricity demand and associated carbon footprint, which may conflict with national climate commitments unless paired with low-carbon power procurement and efficiency measures. For instance, the training of the 176176B BLOOM model consumed $433,196$ kWh total energy and was estimated to emit about $24.69$ tCO2[57]***tCO2e: (tonnes of carbon dioxide equivalent), a standard measure that expresses the total climate impact of greenhouse gas emissions on a common CO2 basis.. Importantly, environmental impacts do not end at training: the inference stage can accumulate substantial footprint at scale. Mistral’s lifecycle assessment [95] reports that by January 2025, Mistral Large 2’s training plus 18 months of usage amounted to $20.4$ ktCO2e, $281,000$ m3 of water consumption, and $660$ kg Sb eqkg Sb eq: (Kilograms of antimony equivalent), a life-cycle assessment metric indicating depletion of scarce mineral resources, normalized to antimony as a reference. of resource depletion. On average, a single $400$-token response (excluding end-user devices) contributes an estimated $1.14$ gCO2e, $45$ mL of water, and $0.16$ mg Sb eq [95]. Under a buying strategy, these impacts are often operationally externalized to the provider’s infrastructure, but governments may still face indirect accountability through procurement standards, public reporting expectations, and alignment with national sustainability targets.
@@ 4.6 National Context Fit Perspective
  • Language: Language coverage is one of the core determinants of the decision especially in multilingual societies. If the primary language is already well supported by global tech companies, buying can deliver strong performance with minimal upfront effort. In contrast, for the purpose of serving minority languages and regional dialects where commercial models may struggle with, the building strategy allows governments to prioritize these low-resource languages using their own specialized datasets. Building can therefore improve inclusive access to public services and help preserve linguistic identity, but it requires sustained effort in data collection, evaluation, and ongoing maintenance.
  • Cultural Fit: The ability to handle historical references, political context, and local social norms is vital for maintaining public trust. Cultural fit involves more than just language; it requires the model to respect religious beliefs, understand local taboos, know local cuisines and follow regional customs. Furthermore, many countries require fine-grained control to account for different laws or values across various states and provinces. In some countries, rich cultural traditions are gradually eroding as younger generations increasingly adopt global cultural norms. Governments may view culturally aligned LLMs as a tool to support cultural continuity, particularly in education, public communication and youth engagement. A locally developed or highly customized model can be programmed with these specific sensitivities in mind, ensuring that its outputs are not only accurate but also culturally appropriate and socially responsible.
  • Legal Interpretations: Can LLMs embed national legal frameworks, case law standards, and procedural rules into its parameters? An LLM should be able to embed national statutes, case law standards, and administrative rules directly into its parameters or retrieval systems. In practice, general-purpose models could default to assumptions from dominant jurisdictions from foreign laws, producing answers that may be misleading, inappropriate, or even unlawful in other countries. When pursuing a buying strategy, governments may therefore need to explicitly negotiate legal disclaimers and domain constraints with model providers to mitigate these risks. On the other hand, building carries risk: without sufficient data and building capability, such systems can produce misleading, hallucinated outputs or even illegal advice [105], wasting resources and undermining trust. A recent example is New York City’s internally developed legal chatbot, which was shut down after producing unreliable advice despite costing over $\mathdollar 600{,}000$ to build [105]. In such cases, combining reliable priopriety models with controlled retrieval (e.g., RAG) may be the safer choice.
@@ 4.7 Evolving Cost-Capability Landscape Perspective
  • Temporal Cost Dynamics: The buy-build decision can shift substantially over time as the underlying cost–capability frontier moves. Training costs may decline rapidly; for example, DeepSeek V3 reports a training cost of roughly \$5.6M, an order-of-magnitude reduction relative to earlier frontier-scale efforts. In parallel, the open-source ecosystem continues to mature, with open models increasingly approaching frontier performance. Compute hardware is also becoming more widely available, while local talent pools can deepen through knowledge transfer and the growing transparency of technical reports and academic publications. Together, these potential trends can reduce the cost and raise the feasibility of building, such that a government may rationally choose buy now and build in the next few years, not because its preferences changed, but because the underlying cost–capability landscape shifted.
  • Capability Debt: LLM development capability is not a function of capital investment alone; it is cumulative, depending on hands-on training experience, institutional know-how, data pipelines and partnerships, and operational lessons learned through iteration and failure. Delaying capability building can therefore create capability debt, raising the cost and difficulty of catching up even as absolute compute costs decline.
  • Option Value of Modest Early Investment: Modest early investments such as small-scale training pilots, sovereign benchmarks, and infrastructure prototyping can preserve future build options without immediate commitment to full-scale training. These efforts accumulate organizational know-how, reduce future transition costs, and enable faster scaling if priorities or constraints change. This option value is often omitted in static cost comparisons but can be strategically material.
  • Re-evaluation Triggers: Governments can make regular re-evaluation by setting clear triggers, such as widening capability gaps, major improvements in open models, vendor price or policy changes, new regulations, or clear progress in domestic talent and infrastructure. Defined triggers and review cycles ensure the strategy keeps pace with the technology landscape rather than changing only in response to crises.